a drag racer starts from rest, accelerates at 4.43 m/s^2. How much time would it require to cover a distance of 200m?

a) 9.5s

b)20.4 s

c)90.3s

d)6.7s
Answer #1

U = 0

a = 4.43 m/s^2

S = 200 m

Use second equation of motion

S = ut + 1/2 at^2

200 = 0 + 0.5 × 4.43 × t^2

t^2 = 90.29

t = 9.5 seconds